OpenStudy (anonymous):

7xy+xy=4 is it linear yes or no. If no tell why?? Help anyone!

OpenStudy (anonymous):

no... just solve for y xy(7+1)=4 xy(8)=4 xy=4/8 y=.5/x

OpenStudy (anonymous):

No, a linear equation's general form is y=m*x + b m & b are just numbers.

OpenStudy (anonymous):

well okay, to add to liza2, .5/x is the same as .5(x^-1) which is not linear.